# Exactly solvable Schrödinger operators related to the hypergeometric equation

### v1

@@ A.3 Degenerate case
$\Big{(}\frac{\alpha+\beta-\mu+1}{2}\Big{)}_{\mu}\Big{(}\frac{\alpha-\beta-\mu+1}{2}\Big{)}_{\mu}$ $F\Big{(}\frac{\alpha+\beta+\mu+1}{2},\frac{\alpha-\beta+\mu+1}{2};1+\mu;z\Big{)}$ (A.17)
$=$ $z^{-\mu}F\Big{(}\frac{\alpha+\beta-\mu+1}{2},\frac{\alpha-\beta-\mu+1}{2};1-\mu;z\Big{)}.$ (A.18)
A.4 Gegenbauer functions
@@
Recall that the Gegenbauer equation is defined by the operator ${\cal G}_{\alpha,\lambda}(w,\partial_{w})$ defined in (2.19), and the two Gegenbauer functions that we use were defined in (2.21) and (2.22). Let us rewrite their definitions using the notation introduced in (A.1).
@@
$S_{\alpha,\pm\lambda}(w)=F_{\alpha,\alpha,\lambda}\left(\frac{1-w}{2}\right).$ (A.19)
It can be easily seen from (A.3) that the solution behaving as $(\frac{w+1}{2})^{-\alpha}$ at $1$ is
$S_{-\alpha,\pm\lambda}(w)=\Big{(}\frac{1+w}{2}\Big{)}^{-\alpha}F_{-\alpha,\alpha,2\lambda}\left(\frac{1-w}{2}\right)$ (A.20)
@@
$Z_{\alpha,\lambda}(w)=(w\pm 1)^{-\frac{1}{2}-\alpha-\lambda}F_{2\lambda,\alpha,\alpha}\left(\frac{2}{1\pm w}\right)$ (A.21)
@@
$\mathbf{S}_{\alpha,\lambda}(w)=\frac{S_{\alpha,\lambda}(w)}{\Gamma(1+\alpha)},\qquad\mathbf{Z}_{\alpha,\lambda}(w)=\frac{Z_{\alpha,\lambda}(w)}{\Gamma(1+\lambda)}.$ (A.22)
We can read their connection formula from the connection formulas of hypergeometric functions (A.13) (A.9), for ${\rm Im}(w)<0$:
$\mathbf{Z}_{\alpha,\lambda}(w)$ $=\frac{\sqrt{\pi}2^{-\alpha+\lambda-\frac{1}{2}}}{\sin(-\pi\alpha)~\Gamma\left(-\alpha+\lambda+\frac{1}{2}\right)}\mathbf{S}_{\alpha,\lambda}(w)+\frac{\sqrt{\pi}2^{\alpha+\lambda-\frac{1}{2}}(w^{2}-1)_{\bullet}^{-\alpha}}{\sin\pi\alpha~\Gamma\left(\alpha+\lambda+\frac{1}{2}\right)}\mathbf{S}_{-\alpha,\lambda}(w),$ (A.23)
$\mathbf{S}_{\alpha,\lambda}(-w)$ $=\frac{\cos\pi\lambda}{\sin(-\pi\alpha)}\mathbf{S}_{\alpha,\lambda}(w)+\frac{\pi~(1-w^{2})^{-\alpha}}{\sin\pi\alpha~\Gamma\left(\frac{1}{2}+\alpha+\lambda\right)\Gamma\left(\frac{1}{2}+\alpha-\lambda\right)}\mathbf{S}_{-\alpha,\lambda}(w).$ (A.24)
@@
$(w^{2}-1)^{\alpha}_{\bullet}:=(w-1)^{\alpha}(w+1)^{\alpha}.$ (A.25)
@@ A.5 Whipple transformation
$\mathbf{Z}_{\alpha,\lambda}(w)$ $=(w^{2}-1)_{\bullet}^{-\frac{1}{4}-\frac{\alpha}{2}-\frac{\lambda}{2}}\mathbf{S}_{\lambda,\alpha}\left(\frac{w}{\left(w^{2}-1\right)_{\bullet}^{\frac{1}{2}}}\right),$ (A.26)
$\mathbf{S}_{\alpha,\lambda}(w)$ $=(w^{2}-1)_{\bullet}^{-\frac{1}{4}-\frac{\alpha}{2}-\frac{\lambda}{2}}\mathbf{Z}_{\lambda,\alpha}\left(\frac{w}{\left(w^{2}-1\right)_{\bullet}^{\frac{1}{2}}}\right),\quad{\rm Re}(w)>0.$ (A.27)
(A.27) is obtain by inverting (A.26) and using the fact that $w\mapsto\frac{w}{(w^{2}-1)^{\frac{1}{2}}_{\bullet}}$ is an involution if and only if ${\rm Re}(w)>0$.
A.6 Half integer case
@@
${\cal G}_{-\frac{1}{2},\lambda}\left(w,\partial_{w}\right)$ $=\partial_{\phi}^{2}+\lambda^{2};$ (A.28)
${\cal G}_{\frac{1}{2},\lambda}\left(w,\partial_{w}\right)$ $=\frac{1}{\sin\phi}\big{(}\partial_{\phi}^{2}+\lambda^{2}\big{)}\sin\phi.$ (A.29)
@@
${\cal G}_{-\frac{1}{2},\lambda}\left(w,\partial_{w}\right)$ $=-\partial_{\theta}^{2}+\lambda^{2};$ (A.30)
${\cal G}_{\frac{1}{2},\lambda}\left(w,\partial_{w}\right)$ $=\frac{1}{\sinh\theta}\big{(}-\partial_{\theta}^{2}+\lambda^{2}\big{)}\sinh\theta.$ (A.31)
(A.28)—(A.31) easily imply the first column of the following identities for Gegenbauer functions. (The second column simply follows from the definitions of Gegenbauer functions).
$\cos\lambda\phi=$ $S_{-\frac{1}{2},\lambda}(\cos\phi)=F\Big{(}\lambda,-\lambda;\frac{1}{2};\sin^{2}\frac{\phi}{2}\Big{)},$ (A.32)
$\frac{\sin\lambda\phi}{\lambda\sin\phi}=$ $S_{\frac{1}{2},\lambda}(\cos\phi)=F\Big{(}1+\lambda,1-\lambda;\frac{3}{2};\sin^{2}\frac{\phi}{2}\Big{)},$ (A.33)
$\cosh\lambda\theta=$ $S_{-\frac{1}{2},\lambda}(\cosh\theta)=F\Big{(}\lambda,-\lambda;\frac{1}{2};\sinh^{2}\frac{\theta}{2}\Big{)},$ (A.34)
$\frac{\sinh\lambda\theta}{\lambda\sinh\theta}=$ $S_{\frac{1}{2},\lambda}(\cosh\theta)=F\Big{(}1+\lambda,1-\lambda;\frac{3}{2};\sinh^{2}\frac{\theta}{2}\Big{)},$ (A.35)
$2^{\lambda}{\rm e}^{-\lambda\theta}=$ $Z_{-\frac{1}{2},\lambda}(\cosh\theta)=\Big{(}2\sinh^{2}\frac{\theta}{2}\Big{)}^{-\lambda}F\Big{(}\lambda,\lambda+\frac{1}{2};1+2\lambda;-\frac{1}{\sinh^{2}\frac{\theta}{2}}\Big{)},$ (A.36)
$\frac{2^{\lambda}{\rm e}^{-\lambda\theta}}{\sinh\theta}=$ $Z_{\frac{1}{2},\lambda}(\cosh\theta)=\Big{(}2\sinh^{2}\frac{\theta}{2}\Big{)}^{-\lambda-1}F\Big{(}\lambda+\frac{1}{2},\lambda+1;1+2\lambda;-\frac{1}{\sinh^{2}\frac{\theta}{2}}\Big{)}.$ (A.37)
@@
$T_{n}(w)=S_{-\frac{1}{2},n}(w),\quad U_{n}(w)=(n+1)S_{\frac{1}{2},n+1}(w).$ (A.38)
Appendix B Closed realizations of 1d Schrödinger operators
B.1 Minimal and maximal realization
@@
$\mathcal{D}(L^{\min})\subset\mathcal{D}(L_{\bullet})\subset\mathcal{D}(L^{\max})$ (B.2)
@@ B.2 Resolvent
${\mathcal{W}}(\Phi_{1},\Phi_{2})(x):=\Phi_{1}(x)\Phi_{2}^{\prime}(x)-\Phi_{1}^{\prime}(x)\Phi_{2}(x).$ (B.3)
@@
${\mathcal{W}}(x):={\mathcal{W}}\big{(}\Psi_{b}(z,\cdot),\Psi_{a}(z,\cdot)\big{)}.$ (B.6)
@@
$R_{\bullet}(z;x,y):=$ $\frac{1}{{\mathcal{W}}(z)}\begin{cases}\Psi_{a}(z,x)\;\!\Psi_{b}(z,y)&\hbox{ if }a<x<y<b,\\ \Psi_{a}(z,y)\;\!\Psi_{b}(z,x)&\hbox{ if }a<y<x<b.\end{cases}$ (B.7)
Note that (B.7) does not depend on the choice of $\Psi_{a}(z,\cdot)$ and $\Psi_{b}(z,\cdot)$.
Suppose that (B.7) defines a bounded operator $R_{\bullet}(z)$, Then $z\not\in\sigma(L_{\bullet})$ and
$\frac{1}{L_{\bullet}-z}=R_{\bullet}(z).$ (B.8)
@@
Appendix C Associated Legendre functions vs. Gegenbauer functions
@@
$\mathcal{L}_{\mu}^{\alpha}:=(1+z^{2})\partial_{z}^{2}-2z\partial_{z}+\mu(\mu+1)-\frac{\alpha^{2}}{1-z^{2}}.$ (C.1)
@@
$(1-w^{2})^{\mp\frac{\alpha}{2}}\mathcal{L}_{\mu}^{\alpha}(1-w^{2})^{\pm\frac{\alpha}{2}}$ (C.2)
$=$ $(1-w^{2})\partial_{w}^{2}-2(\pm\alpha+1)w\partial_{w}+(\mu\mp\alpha)(\mu\pm\alpha+1)={\cal G}_{\pm\alpha,\mu+\frac{1}{2}}.$ (C.3)
@@
$\mathbf{P}^{\alpha}_{\mu}(z)$ $=\left(\frac{z+1}{z-1}\right)^{\frac{\alpha}{2}}\mathbf{F}\left(\mu+1,-\mu;1-\alpha;\frac{1-z}{2}\right)$ (C.4)
$=\frac{2^{\alpha}}{\left(z^{2}-1\right)^{\alpha}_{\bullet}}\mathbf{S}_{-\alpha,\mu+\frac{1}{2}}(z)$ (C.5)
@@
$\mathcal{P}^{\alpha}_{\mu}(z)$ $=\left(\frac{z+1}{1-z}\right)^{\frac{\alpha}{2}}\mathbf{F}\left(\mu+1,-\mu;1-\alpha;\frac{1-z}{2}\right)$ (C.6)
$=\frac{2^{\alpha}}{\left(1-z^{2}\right)^{\alpha}_{\bullet}}\mathbf{S}_{-\alpha,\mu+\frac{1}{2}}(z)$ (C.7)
@@
$\mathbf{Q}^{\alpha}_{\mu}(z)={\rm e}^{{\rm i}\pi\alpha}\frac{\pi^{\frac{1}{2}}\Gamma(\alpha+\mu+1)\left(z^{2}-1\right)^{\frac{\alpha}{2}}}{2^{\mu+1}z^{\alpha+\mu+1}}\mathbf{F}\left(\frac{\alpha+\mu}{2}+1,\frac{\alpha+\mu}{2}+\frac{1}{2};\frac{3}{2}+\mu;z^{-2}\right)$ (C.8)
$={\rm e}^{{\rm i}\pi\alpha}\frac{\pi^{\frac{1}{2}}\Gamma(\alpha+\mu+1)\left(z^{2}-1\right)^{\frac{\alpha}{2}}}{2^{\mu+1}}\mathbf{Z}_{\alpha,\mu+\frac{1}{2}}(z)$ (C.9)
@@
$\mathcal{P}_{\mu}^{\alpha}(\cos r)$ $=\Big{(}\frac{2}{\sin r}\Big{)}^{\frac{1}{2}}\mathcal{P}_{-\alpha,\mu+\frac{1}{2}}^{\rm s}(r),$ (C.10)
${\bf P}_{\mu}^{\alpha}(\cosh r)$ $=\Big{(}\frac{2}{\sinh r}\Big{)}^{\frac{1}{2}}\mathcal{P}_{-\alpha,\mu+\frac{1}{2}}^{\rm h}(r),$ (C.11)
$\mathbf{Q}_{\mu}^{\alpha}(\cosh r)$ $=\frac{1}{(2\sinh r)^{\frac{1}{2}}}{\rm e}^{{\rm i}\pi\alpha}\sqrt{\pi}\Gamma(\alpha+\mu+1)\mathcal{Q}_{-\alpha,\mu+\frac{1}{2}}^{\rm h}(r).$ (C.12)

### v2

@@ C.3 Degenerate case
$\left(\frac{\alpha+\beta-\mu+1}{2}\right)_{\mu}\left(\frac{\alpha-\beta-\mu+1}{2}\right)_{\mu}$ $F\left(\frac{\alpha+\beta+\mu+1}{2},\frac{\alpha-\beta+\mu+1}{2};1+\mu;z\right)$
$=$ $z^{-\mu}F\left(\frac{\alpha+\beta-\mu+1}{2},\frac{\alpha-\beta-\mu+1}{2};1-\mu;z\right).$ (C.17)
C.4 Gegenbauer functions
@@
Recall that the Gegenbauer equation is defined by the operator ${\cal G}_{\alpha,\lambda}(w,\partial_{w})$ defined in (2.7), and the two Gegenbauer functions that we use were defined in (3.1) and (2.10). Let us rewrite their definitions using the notation introduced in (C.1).
@@
$S_{\alpha,\pm\lambda}(w)=F_{\alpha,\alpha,\lambda}\left(\frac{1-w}{2}\right).$ (C.18)
It can be easily seen from (C.3) that the solution behaving as $(\frac{w+1}{2})^{-\alpha}$ at $1$ is
$S_{-\alpha,\pm\lambda}(w)=\Big(\frac{1+w}{2}\Big)^{-\alpha}F_{-\alpha,\alpha,2\lambda}\left(\frac{1-w}{2}\right)$ (C.19)
@@
$Z_{\alpha,\lambda}(w)=(w\pm 1)^{-\frac{1}{2}-\alpha-\lambda}F_{2\lambda,\alpha,\alpha}\left(\frac{2}{1\pm w}\right)$ (C.20)
@@
$\mathbf{S}_{\alpha,\lambda}(w)=\frac{S_{\alpha,\lambda}(w)}{\Gamma(1+\alpha)},\qquad\mathbf{Z}_{\alpha,\lambda}(w)=\frac{Z_{\alpha,\lambda}(w)}{\Gamma(1+\lambda)}.$ (C.21)
We can read their connection formula from the connection formulas of hypergeometric functions (C.13) (C.9), for ${\rm Im}(w)<0$:
$\mathbf{Z}_{\alpha,\lambda}(w)$ $=\frac{\sqrt{\pi}2^{-\alpha+\lambda-\frac{1}{2}}}{\sin(-\pi\alpha)~\Gamma\left(-\alpha+\lambda+\frac{1}{2}\right)}\mathbf{S}_{\alpha,\lambda}(w)+\frac{\sqrt{\pi}2^{\alpha+\lambda-\frac{1}{2}}(w^{2}-1)_{\bullet}^{-\alpha}}{\sin\pi\alpha~\Gamma\left(\alpha+\lambda+\frac{1}{2}\right)}\mathbf{S}_{-\alpha,\lambda}(w),$ (C.22)
$\mathbf{S}_{\alpha,\lambda}(-w)$ $=\frac{\cos\pi\lambda}{\sin(-\pi\alpha)}\mathbf{S}_{\alpha,\lambda}(w)+\frac{\pi~(1-w^{2})^{-\alpha}}{\sin\pi\alpha~\Gamma\left(\frac{1}{2}+\alpha+\lambda\right)\Gamma\left(\frac{1}{2}+\alpha-\lambda\right)}\mathbf{S}_{-\alpha,\lambda}(w).$ (C.23)
@@
$(w^{2}-1)^{\alpha}_{\bullet}:=(w-1)^{\alpha}(w+1)^{\alpha}.$ (C.24)
@@ C.5 Whipple transformation
$\mathbf{Z}_{\alpha,\lambda}(w)$ $=(w^{2}-1)_{\bullet}^{-\frac{1}{4}-\frac{\alpha}{2}-\frac{\lambda}{2}}\mathbf{S}_{\lambda,\alpha}\left(\frac{w}{\left(w^{2}-1\right)_{\bullet}^{\frac{1}{2}}}\right),$ (C.25)
$\mathbf{S}_{\alpha,\lambda}(w)$ $=(w^{2}-1)_{\bullet}^{-\frac{1}{4}-\frac{\alpha}{2}-\frac{\lambda}{2}}\mathbf{Z}_{\lambda,\alpha}\left(\frac{w}{\left(w^{2}-1\right)_{\bullet}^{\frac{1}{2}}}\right),\quad{\rm Re}(w)>0.$ (C.26)
(C.26) is obtain by inverting (C.25) and using the fact that $w\mapsto\frac{w}{(w^{2}-1)^{\frac{1}{2}}_{\bullet}}$ is an involution if and only if ${\rm Re}(w)>0$.
C.6 Half integer case
@@
${\cal G}_{-\frac{1}{2},\lambda}\left(w,\partial_{w}\right)$ $=\partial_{\phi}^{2}+\lambda^{2};$ (C.27)
${\cal G}_{\frac{1}{2},\lambda}\left(w,\partial_{w}\right)$ $=\frac{1}{\sin\phi}\big(\partial_{\phi}^{2}+\lambda^{2}\big)\sin\phi.$ (C.28)
@@
${\cal G}_{-\frac{1}{2},\lambda}\left(w,\partial_{w}\right)$ $=-\partial_{\theta}^{2}+\lambda^{2};$ (C.29)
${\cal G}_{\frac{1}{2},\lambda}\left(w,\partial_{w}\right)$ $=\frac{1}{\sinh\theta}\big(-\partial_{\theta}^{2}+\lambda^{2}\big)\sinh\theta.$ (C.30)
(C.27)—(C.30) easily imply the first column of the following identities for Gegenbauer functions. (The second column simply follows from the definitions of Gegenbauer functions).
$\cos\lambda\phi=$ $S_{-\frac{1}{2},\lambda}(\cos\phi)=F\Big(\lambda,-\lambda;\frac{1}{2};\sin^{2}\frac{\phi}{2}\Big),$ (C.31)
$\frac{\sin\lambda\phi}{\lambda\sin\phi}=$ $S_{\frac{1}{2},\lambda}(\cos\phi)=F\Big(1+\lambda,1-\lambda;\frac{3}{2};\sin^{2}\frac{\phi}{2}\Big),$ (C.32)
$\cosh\lambda\theta=$ $S_{-\frac{1}{2},\lambda}(\cosh\theta)=F\Big(\lambda,-\lambda;\frac{1}{2};\sinh^{2}\frac{\theta}{2}\Big),$ (C.33)
$\frac{\sinh\lambda\theta}{\lambda\sinh\theta}=$ $S_{\frac{1}{2},\lambda}(\cosh\theta)=F\Big(1+\lambda,1-\lambda;\frac{3}{2};\sinh^{2}\frac{\theta}{2}\Big),$ (C.34)
$2^{\lambda}{\rm e}^{-\lambda\theta}=$ $Z_{-\frac{1}{2},\lambda}(\cosh\theta)=\Big(2\sinh^{2}\frac{\theta}{2}\Big)^{-\lambda}F\Big(\lambda,\lambda+\frac{1}{2};1+2\lambda;-\frac{1}{\sinh^{2}\frac{\theta}{2}}\Big),$ (C.35)
$\frac{2^{\lambda}{\rm e}^{-\lambda\theta}}{\sinh\theta}=$ $Z_{\frac{1}{2},\lambda}(\cosh\theta)=\Big(2\sinh^{2}\frac{\theta}{2}\Big)^{-\lambda-1}F\Big(\lambda+\frac{1}{2},\lambda+1;1+2\lambda;-\frac{1}{\sinh^{2}\frac{\theta}{2}}\Big).$ (C.36)
@@
$T_{n}(w)=S_{-\frac{1}{2},n}(w),\quad U_{n}(w)=(n+1)S_{\frac{1}{2},n+1}(w).$ (C.37)
Appendix D Closed realizations of 1d Schrödinger operators
D.1 Minimal and maximal realization
@@
$\mathcal{D}(L^{\min})\subset\mathcal{D}(L_{\bullet})\subset\mathcal{D}(L^{\max})$ (D.2)
@@ D.2 Resolvent
${\mathcal{W}}(\Phi_{1},\Phi_{2})(x):=\Phi_{1}(x)\Phi_{2}^{\prime}(x)-\Phi_{1}^{\prime}(x)\Phi_{2}(x).$ (D.3)
@@
${\mathcal{W}}(x):={\mathcal{W}}\big(\Psi_{b}(z,\cdot),\Psi_{a}(z,\cdot)\big).$ (D.6)
@@
$R_{\bullet}(z;x,y):=$ $\frac{1}{{\mathcal{W}}(z)}\begin{cases}\Psi_{a}(z,x)\;\!\Psi_{b}(z,y)&\hbox{ if }a<x<y<b,\\ \Psi_{a}(z,y)\;\!\Psi_{b}(z,x)&\hbox{ if }a<y<x<b.\end{cases}$ (D.7)
Note that (D.7) does not depend on the choice of $\Psi_{a}(z,\cdot)$ and $\Psi_{b}(z,\cdot)$.
Suppose that (D.7) defines a bounded operator, which we denote $R_{\bullet}(z)$. Then
$z\not\in\sigma(L_{\bullet})\quad\text{and }\quad\frac{1}{L_{\bullet}-z}=R_{\bullet}(z).$ (D.8)
@@
Above we described the path from $L_{\bullet}$ to $R_{\bullet}(z)$. In our constructions we will use the reverse direction, from $R_{\bullet}(z)$ to $L_{\bullet}$, which is described in the following theorem.
Theorem D.1.
Suppose that $z\in{\mathbb{C}}$ and let $\Psi_{a}(z,\cdot),\Psi_{b}(z,\cdot)$ be functions in $AC^{1}]a,b[$ solving the eigenvalue equations (D.4), resp. (D.5), such that $\Psi_{a}(z,\cdot)$ is in $L^{2}$ near $a$ and $\Psi_{b}(z,\cdot)$ is in $L^{2}$ near $b$. Define ${\mathcal{W}}(x)$ by (D.6) and $R_{\bullet}(z;x,y)$ by (D.7). Suppose that (D.7) defines a bounded operator on $L^{2}]a,b[$, which we denote $R_{\bullet}(z)$. Then there exists a unique operator closed $L_{\bullet}$ that satisfies (D.8).
Appendix E Associated Legendre functions vs. Gegenbauer functions
@@
$\mathcal{L}_{\mu}^{\alpha}:=(1+z^{2})\partial_{z}^{2}-2z\partial_{z}+\mu(\mu+1)-\frac{\alpha^{2}}{1-z^{2}}.$ (E.1)
@@
$(1-w^{2})^{\mp\frac{\alpha}{2}}\mathcal{L}_{\mu}^{\alpha}(1-w^{2})^{\pm\frac{\alpha}{2}}$
$=$ $(1-w^{2})\partial_{w}^{2}-2(\pm\alpha+1)w\partial_{w}+(\mu\mp\alpha)(\mu\pm\alpha+1)={\cal G}_{\pm\alpha,\mu+\frac{1}{2}}.$ (E.2)
@@
$\mathbf{P}^{\alpha}_{\mu}(z)$ $=\left(\frac{z+1}{z-1}\right)^{\frac{\alpha}{2}}\mathbf{F}\left(\mu+1,-\mu;1-\alpha;\frac{1-z}{2}\right)$
$=\frac{2^{\alpha}}{\left(z^{2}-1\right)^{\alpha}_{\bullet}}\mathbf{S}_{-\alpha,\mu+\frac{1}{2}}(z)$ (E.3)
@@
$\mathcal{P}^{\alpha}_{\mu}(z)$ $=\left(\frac{z+1}{1-z}\right)^{\frac{\alpha}{2}}\mathbf{F}\left(\mu+1,-\mu;1-\alpha;\frac{1-z}{2}\right)$
$=\frac{2^{\alpha}}{\left(1-z^{2}\right)^{\alpha}_{\bullet}}\mathbf{S}_{-\alpha,\mu+\frac{1}{2}}(z)$ (E.4)
@@
$\mathbf{Q}^{\alpha}_{\mu}(z)$ $={\rm e}^{{\rm i}\pi\alpha}\frac{\pi^{\frac{1}{2}}\Gamma(\alpha+\mu+1)\left(z^{2}-1\right)^{\frac{\alpha}{2}}}{2^{\mu+1}z^{\alpha+\mu+1}}\mathbf{F}\left(\frac{\alpha+\mu}{2}+1,\frac{\alpha+\mu}{2}+\frac{1}{2};\frac{3}{2}+\mu;z^{-2}\right)$
$={\rm e}^{{\rm i}\pi\alpha}\frac{\pi^{\frac{1}{2}}\Gamma(\alpha+\mu+1)\left(z^{2}-1\right)^{\frac{\alpha}{2}}}{2^{\mu+1}}\mathbf{Z}_{\alpha,\mu+\frac{1}{2}}(z).$ (E.5)
@@
$\mathcal{P}_{\mu}^{\alpha}(\cos r)$ $=\Big(\frac{2}{\sin r}\Big)^{\frac{1}{2}}\mathcal{P}_{-\alpha,\mu+\frac{1}{2}}^{\rm s}(r),$ (E.6)
${\bf P}_{\mu}^{\alpha}(\cosh r)$ $=\Big(\frac{2}{\sinh r}\Big)^{\frac{1}{2}}\mathcal{P}_{-\alpha,\mu+\frac{1}{2}}^{\rm h}(r),$ (E.7)
$\mathbf{Q}_{\mu}^{\alpha}(\cosh r)$ $=\frac{1}{(2\sinh r)^{\frac{1}{2}}}{\rm e}^{{\rm i}\pi\alpha}\sqrt{\pi}\Gamma(\alpha+\mu+1)\mathcal{Q}_{-\alpha,\mu+\frac{1}{2}}^{\rm h}(r).$ (E.8)
Conflict of interest. This manuscript has no conflict of interest.